# Magnetically charged black-bounce solution via nonlinear electrodynamics in a k-essence theory

### v1

@@ V.1 Energy conditions for $n=1/3$
As the energy conditions depend directly on the scalar field, these components of the stress-energy tensor are to the field configuration where $n=1/3$. For the region $A>0$, the components are defined as:
@@
${p}^{\phi}_{2}$ $=$ ${p}^{\phi}_{3}=\frac{3A{\Sigma}^{\prime\prime}}{\Sigma}-V(x),\quad\quad{p}^{EM}_{2}={p}^{EM}_{3}=-\frac{L(x)}{2}+\frac{q^{2}_{m}{L_{f}(x)}}{2\Sigma^{4}}.$ (54)
@@
${p}^{\phi}_{2}$ $=$ ${p}^{\phi}_{3}=\frac{3A{\Sigma}^{\prime\prime}}{\Sigma}-V(x),\quad\quad{p}^{EM}_{2}={p}^{EM}_{3}=-\frac{L(x)}{2}+\frac{q^{2}_{m}{L_{f}(x)}}{2\Sigma^{4}}.$ (57)
@@
The electromagnetic part within the horizon is the same expressions obtained above in Eqs. (64-69).
@@ V.2 Energy conditions for $n=1/5$
${p}^{\phi}_{2}$ $=$ ${p}^{\phi}_{3}=\frac{5A{\Sigma}^{\prime\prime}}{\Sigma}-V(x),\quad\quad{p}^{EM}_{2}={p}^{EM}_{3}=-\frac{L(x)}{2}+\frac{q^{2}_{m}{L_{f}(x)}}{2\Sigma^{4}}.$ (78)
@@
${p}^{\phi}_{2}$ $=$ ${p}^{\phi}_{3}=\frac{5A{\Sigma}^{\prime\prime}}{\Sigma}-V(x),\quad\quad{p}^{EM}_{2}={p}^{EM}_{3}=-\frac{L(x)}{2}+\frac{q^{2}_{m}{L_{f}(x)}}{2\Sigma^{4}}.$ (81)
@@
The internal energy conditions at the horizon $A<0$ for this field configuration are given by
@@
The electromagnetic part within the horizon is the same expressions obtained above in Eqs. (88-93).
@@ VI.1 Simpson-Visser model
The energy conditions for the scalar field and electromagnetism in the region external to the event horizon $A>0$ and considering the $n=1/3$ configuration, are set out explicitly below, where the quantities used will be used Eqs. (20-23)
@@
Analyzing the energy conditions for the scalar field outside the event horizon ($A>0$), we observe that the null energy condition $NEC^{\phi}_{1}$ Eq. (58) is violated, implying a violation of the dominant energy condition $DEC^{\phi}_{1}$. All other energy conditions for the scalar field are violated except for $NEC^{\phi}_{2}$ Eq. (59). Regarding the energy conditions of the electromagnetic part, all of them are satisfied everywhere, except for the dominant energy condition $DEC^{EM}_{2}$ Eq. (104).
@@
Note that the electromagnetic energy conditions within the event horizon are the same as those defined for the outer region Eqs. (103-105).
@@
VI.2 Bardeen type model
Replacing the metric functions in the energy conditions, in the region where $A(x)>0$, Eqs. (82-87), we have:
@@ VI.2 Bardeen type model
For the electromagnetic part, the main null energy condition $NEC^{EM}_{1}$ Eq. (64) is satisfied within the horizon, as well as the dominant energy conditions $DEC^{EM}_{1}$ Eq. (112) and $DEC^{EM}_{3}$ Eq. (113). The secondary null energy condition $NEC^{EM}_{2}$ Eq. (111) is violated for throat radii between $-\sqrt{2/13}q_{m}<x<\sqrt{2/13}q_{m}$. Likewise, the dominant energy condition $DEC^{EM}_{2}$ Eq. (112) is violated for throat radii between $-\sqrt{134/91}q_{m}>x>\sqrt{134/91}q_{m}$. Finally, the strong energy condition $SEC^{EM}_{3}$ Eq. (111) is violated for throat radii between $-\sqrt{34/91}q_{m}<x<\sqrt{34/91}q_{m}$.
@@ VII Conclusion
In the present work, we start from an action that describes the k-essence theory with a scalar field assuming a power form. This theory has been used to investigate black-bounce solutions [45] and generalizations for different scalar field configurations [46]. We extend this framework by introducing non-linear electrodynamics [52] to explore possible charged black-bounce solutions for scalar field strengths differing from the canonical case $n=1$. Specifically, we construct magnetic solutions such that the throat of the wormhole coincides with the magnetic charge $a=q_{m}$.
@@
From the equations of motion, Eqs. (15-18), we can observe that the electromagnetic function $L_{f}$, contained in the equation Eq. (16), does not depend on the form of the scalar field but rather only of the metric functions $A(x)$ and $\Sigma(x)$. Which leads to the conclusion that this function is the same as obtained for the canonical scalar field depending only on the chosen model.
@@
We analytically derive all the functions involved for the Simpson-Visser models (Section III) and for a Bardeen-type solution (Section IV) for k-essence configurations with $n=1/3$ and $n=1/5$. We can verify that for each model of specific form the electromagnetic functions $L_{f}(x)$ and $L(x)$ do not change with the variation in the power of the k-essence field. In fact, these functions are the same as those obtained in the canonical case and investigated in [15, 52]. This behavior implies that the modifications in the scalar field due to k-essence are counterbalanced by the potential, keeping $L_{f}(x)$ and $L(x)$ unchanged. This behavior can be extended to the other powers of the phantom scalar field $n=1/3,1/5,1/7,\dots$.
@@
We graphically represent the scalar field, potential, and electromagnetic functions for each of the models studied. Qualitatively, the electromagnetic functions $L_{f}(x)$ and $L(x)$ for both the Simpson-Visser model (Section 1) and the Bardeen model (Section 3) exhibit similar behavior. Regarding the behavior of the scalar field for both models in this configuration, it tends to invert its sign asymptotically $x\to{\pm\infty}$ for some charge values, as illustrated in Eqs. (27) and (40). The potential for both models and configurations tends to behave similar to a potential barrier for regions outside the horizon and creates minima as it becomes more internal to the horizon.
@@
Finally, we analyzed the energy conditions for the scalar and electromagnetic fields for each of the models investigated. As previously observed in [46], the energy conditions for the scalar field do not change with the power of the k-essence, so it suffices to analyze only the case of the lower power $n=1/3$. Thus, in the Simpson-visser model, VI.1, the null energy condition $NEC^{\phi}_{1}$ and its associated dominant energy condition $DEC^{\phi}_{1}$ are violated outside the horizon, $A>0$. As well as all other energy conditions, except for $NEC^{\phi}_{2}$. For within the event horizon, $A<0$, all energy conditions are violated, Eqs. (106-107). Regarding the analysis of electromagnetic energy conditions for the Simpson-Visser model, all conditions are satisfied everywhere, except $DEC^{EM}_{2}$, Eq. (104).
@@
In the analysis of the energy conditions of the second model VI.2, the energy conditions $NEC^{\phi}_{1}$, $DEC^{\phi}_{1}$, $DEC^{\phi}_{2}$, and $DEC^{\phi}_{3}$ are violated where $A>0$. The null energy condition $NEC^{\phi}_{2}$ is satisfied and $SEC^{\phi}_{3}$, Eq. (108), is conditionally violated. Regarding the electromagnetic part, the energy conditions $NEC^{EM}_{1}$, $DEC^{EM}_{1}$ and $DEC^{EM}_{3}$ are satisfied and the others are conditionally violated Eqs. (111-113). In regions where $A<0$, the null energy conditions $NEC^{\phi}_{1}$ and $NEC^{\phi}_{2}$ are violated Eq. (114), leading to the violation of the dominant energy conditions $DEC^{\phi}_{1}$, $DEC^{\phi}_{2}$. The energy density is violated, but $SEC^{\phi}_{3}$, Eq. (116), is satisfied.

### v2

@@ V.1 Energy conditions for $n=1/3$
As the energy conditions depend directly on the scalar field, we will consider these components of the stress-energy tensor to the field configuration where $n=1/3$. For the region $A>0$, the components are defined as:
@@
${p}^{\phi}_{2}$ $=$ $\frac{3A{\Sigma}^{\prime\prime}}{\Sigma}-V(x),\quad\quad{p}^{EM}_{2}=-\frac{L(x)}{2}+\frac{q^{2}_{m}{L_{f}(x)}}{2\Sigma^{4}}.$ (56)
@@
${p}^{\phi}_{2}$ $=$ $\frac{3A{\Sigma}^{\prime\prime}}{\Sigma}-V(x),\quad\quad{p}^{EM}_{2}=-\frac{L(x)}{2}+\frac{q^{2}_{m}{L_{f}(x)}}{2\Sigma^{4}}.$ (59)
@@
The electromagnetic part within the horizon is the same expressions obtained above in Eqs. (66-71).
@@ V.2 Energy conditions for $n=1/5$
${p}^{\phi}_{2}$ $=$ $\frac{5A{\Sigma}^{\prime\prime}}{\Sigma}-V(x),\quad\quad{p}^{EM}_{2}=-\frac{L(x)}{2}+\frac{q^{2}_{m}{L_{f}(x)}}{2\Sigma^{4}}.$ (80)
@@
${p}^{\phi}_{2}$ $=$ $\frac{5A{\Sigma}^{\prime\prime}}{\Sigma}-V(x),\quad\quad{p}^{EM}_{2}=-\frac{L(x)}{2}+\frac{q^{2}_{m}{L_{f}(x)}}{2\Sigma^{4}}.$ (83)
@@
The energy conditions where $A<0$ for this field configuration are given by
@@
The electromagnetic parts within the horizon are the same expressions as obtained above in Eqs. (90-95).
@@ VI.1 Simpson-Visser model
The energy conditions for the scalar field and electromagnetism, in the region external to the event horizon, $A>0$, and considering the $n=1/3$ configuration, are set out explicitly below, where the quantities Eqs. (20-23) will be used.
@@
Analyzing the energy conditions for the scalar field outside the event horizon ($A>0$), we observe that the null energy condition $NEC^{\phi}_{1}$ Eq. (60) is violated, which implies a violation of the dominant energy condition $DEC^{\phi}_{1}$. All other energy conditions for the scalar field are violated except for $NEC^{\phi}_{2}$ Eq. (61). Regarding the energy conditions of the electromagnetic part, they are satisfied everywhere, except for the dominant energy condition $DEC^{EM}_{2}$ Eq. (106).
@@
Note that the electromagnetic energy conditions within the event horizon are the same as those defined for the outer region Eqs. (105-107).
@@
VI.2 Bardeen-type model
Replacing the metric functions in the energy conditions, Eqs. (84-89), in the region where $A(x)>0$, we have:
@@ VI.2 Bardeen-type model
For the electromagnetic part, the main null energy condition $NEC^{EM}_{1}$ Eq. (66) is satisfied within the horizon, as well as the dominant energy conditions $DEC^{EM}_{1}$ Eq. (114) and $DEC^{EM}_{3}$ Eq. (115). The secondary null energy condition $NEC^{EM}_{2}$ Eq. (113) is violated for the throat radii between $-\sqrt{2/13}q_{m}<x<\sqrt{2/13}q_{m}$. Likewise, the dominant energy condition $DEC^{EM}_{2}$ Eq. (114) is violated for the throat radii between $-\sqrt{134/91}q_{m}>x>\sqrt{134/91}q_{m}$. Finally, the strong energy condition $SEC^{EM}_{3}$ Eq. (113) is violated for the throat radii between $-\sqrt{34/91}q_{m}<x<\sqrt{34/91}q_{m}$.
@@ VII Conclusion
In the present work, we start from an action that describes the k-essence theory with a scalar field, assuming a power form. This theory has been used to investigate black-bounce solutions [45] and generalizations for different scalar field configurations [46]. We extend this framework by introducing NED [52] to explore possible charged black-bounce solutions for scalar field strengths that differ from the canonical case $n=1$. Specifically, we construct magnetic solutions such that the throat of the wormhole coincides with the magnetic charge $a=q_{m}$.
@@
From the equations of motion, Eqs. (15-18), we can observe that the electromagnetic function $L_{f}$, contained in the equation Eq. (16), does not depend on the form of the scalar field, but rather only on the metric functions $A(x)$ and $\Sigma(x)$. This leads to the conclusion that this function is the same as obtained for the canonical scalar field depending only on the chosen model.
@@
We analytically derive all the functions involved for the Simpson-Visser model (Section III) and for a Bardeen-type solution (Section IV) for k-essence configurations with $n=1/3$ and $n=1/5$. We can verify that for each model of a specific form, the electromagnetic functions $L_{f}(x)$ and $L(x)$ do not change with variation in the power of the k-essence field. In fact, these functions are the same as those obtained in the canonical case and investigated in [15, 52]. This behavior implies that the modifications in the scalar field due to k-essence are counterbalanced by the potential, keeping $L_{f}(x)$ and $L(x)$ unchanged. This behavior can be extended to the other powers of the phantom scalar field $n=1/3,1/5,1/7,\dots$.
@@
We graphically represent the scalar field, potential, and electromagnetic functions for each of the models studied. Qualitatively, the electromagnetic functions $L_{f}(x)$ and $L(x)$ for both the Simpson-Visser model (Fig. 1) and the Bardeen-type model (Fig. 3) exhibit similar behavior. Regarding the behavior of the scalar field for both models in this configuration, it tends to invert its sign asymptotically $x\to{\pm\infty}$ for some charge values, as illustrated in Eqs. (27) and (41). The potential for both models and configurations tends to behave similar to a potential barrier for regions outside the horizon and creates minima as it becomes more internal to the horizon.
@@
Finally, we analyzed the energy conditions for the scalar and electromagnetic fields for each of the models investigated. As previously observed in [46], the energy conditions for the scalar field do not change with the power of the k-essence, so it suffices to analyze only the case of the lower power $n=1/3$. Thus, in the Simpson-visser model, Section VI.1, the null energy condition $NEC^{\phi}_{1}$ and its associated dominant energy condition $DEC^{\phi}_{1}$ are violated outside the horizon, $A>0$. As well as all other energy conditions, except for $NEC^{\phi}_{2}$. For within the event horizon, $A<0$, all energy conditions are violated, Eqs. (108-109). Regarding the analysis of electromagnetic energy conditions for the Simpson-Visser model, all conditions are satisfied everywhere, except $DEC^{EM}_{2}$, Eq. (106).
@@
In the analysis of the energy conditions of the second model, Section VI.2, the energy conditions $NEC^{\phi}_{1}$, $DEC^{\phi}_{1}$, $DEC^{\phi}_{2}$, and $DEC^{\phi}_{3}$ are violated where $A>0$. The null energy condition $NEC^{\phi}_{2}$ is satisfied and $SEC^{\phi}_{3}$, Eq. (110), is conditionally violated. Regarding the electromagnetic part, the energy conditions $NEC^{EM}_{1}$, $DEC^{EM}_{1}$, and $DEC^{EM}_{3}$ are satisfied, and the others are conditionally violated Eqs. (113-115). In regions where $A<0$, the null energy conditions $NEC^{\phi}_{1}$ and $NEC^{\phi}_{2}$ are violated, Eq. (116), leading to the violation of the dominant energy conditions $DEC^{\phi}_{1}$, $DEC^{\phi}_{2}$. The energy density is violated, but $SEC^{\phi}_{3}$, Eq. (118), is satisfied.